# Approximate Degrees of Multisymmetric Properties with Application to Quantum Claw Detection

### v1

@@ 1.4 Technical Outline
If $h$ is one-to-one, then $h|_{S}$ and $h|_{T}$ obviously have no collision for every $S$ and $T$, where $h|_{S}$ and $h|_{T}$ are the restriction of $h$ to the the subsets $S$ and $T$, respectively, of the domain. Thus, $P$ is close to 00. If $h$ is two-to-one, $h|_{S}$ is injective with high probability and, in this case, $T$ intersects the preimage of $h|_{S}(S)$ with high probability. This follows from the birthday paradox, since $|S|=F<<\sqrt{M}=\sqrt{FG}$ and $|T|=G>>\sqrt{M}$. Therefore, $P$ is far from 00. By scaling and shifting $P$, we have a polynomial that $O(1)$-approximates the collision property.
As an application of Theorem 4, we obtain Theorem 5 by generalizing the reduction from the claw problem for two functions with the same domain size [7] to the reduction from the problem for two functions with the different domain size: $f\colon[F]\to[F+G]$ and $g\colon[G]\to[F+G]$ with $F\leq G\leq F^{2}$ (the case of $G>F^{2}$ is already covered by Theorem 4). More precisely, we reduce the composition of a certain problem (called $\mathsf{pSearch}$ [11]) and the claw problem with the function range stated in Theorem 4 to the claw problem in question, and then use the composition theorem of the general adversary bound [18]. That is, for every $k,F,G\in\mathbb{N}$ with $k|F$ and $G|(F/k)$, the composition $\mathsf{claw}_{(F,G)\to M+2}$ is reducible from the problem $\mathsf{claw}_{(k,kG/F)\to M}\circ\mathsf{pSearch}_{F/k\to M}$, where $\mathsf{claw}_{(a,b)\to c}$ denotes the decision problem of detecting the existence of a claw for given functions $(\phi,\psi)\in[c]^{[a]}\times[c]^{[b]}$. Since $\mathsf{pSearch}_{N\to M}$ is an variant of the unstructured search over $N$ elements, the quantum query complexity of $\mathsf{pSearch}_{F/k\to M}$ is $\Omega(\sqrt{F/k})$. By setting $M=k+kG/F$, Theorem 4 provides a lower bound on the quantum query complexity of $\mathsf{claw}_{(k,kG/F)\to M}$ as
@@
By applying the compositon theorem of the general adversary bound, we have
@@ 1.5 Organization
Sec. missing  2 provides notations and necessary notions. Sec. missing  3 proves Theorem 1 and Theorem 2. Sec. missing  4 proves Theorem 3 and Theorem 4. Sec. missing  5 proves Theorem 5.
@@ 2.1 Notations
[…]
@@ 2.2 Polynomial Method for Quantum Lower bound
The polynomial method was introduced by Beals et al. [9] to obtain lower bounds on the number of quantum queries required to compute a Boolean function on an input bit-string given as an oracle. It is not difficult to generalize the method to the case where the input is a string of alphabets [1]. In the following, we consider the case where the input consists of two alphabet-strings corresponding to two functions. But it is not difficult to see similar facts hold for three or more functions.
The following lemma says that, after $q$ quantum queries, any quantum algorithm is in the state where the amplitude of each computational basis state is represented by a polynomial of degree at most $q$ in $x_{ij}$ and $y_{kj}$.
@@
where $+$ is the addition modulo $M$, or alternatively,
@@
where $\oplus$ is bitwise XOR in binary expression. At any step, the the state of the algorithm is of the form
@@
Let $F,G,M$ be in $\mathbb{N}$, and let $\Phi$ be a partial Boolean function over the set of pairs of functions in $\mathcal{D}\subseteq[M]^{[F]}\times[M]^{[G]}$. For given oracles $O_{f}$ and $O_{g}$ for functions $(f,g)\in\mathcal{D}$, suppose that there exists a quantum algorithm that computes $\Phi(f,g)$ with error $\epsilon$ using $q$ queries. Then, there is a polynomial $P$ of degree at most $2q$ in variables $x_{11},\dots,x_{FM},y_{11},\dots,y_{GM}$ such that $P$ $\epsilon$-approximates $\Phi$, where $x_{ij}$ and $y_{kj}$ are the variables defined for $f$ and $g$, respectively, in Sec. missing  2.2.
@@
For a partial Boolean function $\Phi$ and $20<\epsilon<\delta<1/$ 2, it holds that
@@ 2.3 Multisymmetric Properties
The definition of bisymmetric properties can be generalized to the three or more function case in a straightforward way. Let $\Phi\colon\mathcal{D}\to\{0,1\}$ be a partial Boolean function over the set of $k$-tuples of functions in $\mathcal{D}\subseteq\mathcal{F}_{1}\times\cdots\times\mathcal{F}_{k}$, where $\mathcal{F}_{\ell}\ (\ell\in[k])$ denotes $[M]^{[F_{\ell}]}$ for $F_{\ell},M\in\mathbb{Z}_{+}$. We say that $\Phi$ is $k$-symmetric, if $\Phi(f_{1},\dots,f_{k})$ is invariant and $\mathcal{D}$ is closed under the permutations over $\mathrm{dom}(f_{\ell})=[F_{\ell}]$ for each $\ell\in[k]$, respectively, and the permutation over the common range $[M]$ of $f_{\ell}$’s, i.e., $\left(\sigma f_{1}\pi_{F_{1}},\dots,\sigma f_{k}\pi_{F_{k}}\right)\in\mathcal{D}$ if $(f_{1},\dots,f_{k})\in\mathcal{D}$, and
@@
for every permutation $\pi_{F_{\ell}}$ over $[F_{\ell}]$ for $\ell\in[k]$ and every permutation $\sigma$ over $[M]$, respectively.
@@ 3.1 Bisymmetric Properties
We first show that there are two polynomials in different sets of variables of the same degree such that both polynomials approximates $\Phi$.
@@
For $F,G,M\in\mathbb{N}$ and $\mathcal{D}\subseteq[M]^{[F]}\times[M]^{[G]}$, let $\Phi:\mathcal{D}\to\{0,1\}$ be a bisymmetric partial Boolean function over the set of functions $(f,g)\in\mathcal{D}$. Define $FM$ variables $x_{ij}$ and $GM$ variables $y_{kj}$ for functions $f$ and $g$ as in Sec. missing  2.2. In addition, define $z_{j}$ as the number of preimages of $j$ via $f$, i.e., $z_{j}=\left|f^{-1}\left(j\right)\right|$, and define $w_{j}$ as the number of preimages of $j$ via $g$, i.e., $w_{j}=\left|g^{-1}\left(j\right)\right|$. Then, the following are equivalent.
@@
Next, we show that the degree of a polynomial $Q$ shown in Sec. missing  3.1 does not change, even if we restrict the common range of input functions of $\Phi$ to a small subset.
@@
For $F,G,M\in\mathbb{N}$ such that $M\geq F+G$ and $\mathcal{D}\subseteq[M]^{[F]}\times[M]^{[G]}$, let $\Phi\colon\mathcal{D}\to\{0,1\}$ be a bisymmetric partial Boolean function over the set of function pairs $(f,g)\in\mathcal{D}$. In addition, for every $M^{\prime}\in\mathbb{N}$ such that $F+G\leq M^{\prime}<M$, let $\Phi^{\prime}\colon\mathcal{D}\cap\left([M^{\prime}]^{[F]}\times[M^{\prime}]^{[G]}\right)\to\{0,1\}$ be the restriction of $\Phi$ to the domain $[M^{\prime}]^{[F]}\times[M^{\prime}]^{[G]}$, i.e., the bisymmetric Boolean function over the set of function pairs $(f^{\prime},g^{\prime})\in\mathcal{D}\cap\left([M^{\prime}]^{[F]}\times[M^{\prime}]^{[G]}\right)$ such that $\Phi^{\prime}(f^{\prime},g^{\prime})=\Phi(f,g)$ for all functions $f^{\prime},g^{\prime},f,g$ such that $f^{\prime}(x)=f(x)$ and $g^{\prime}(y)=g(y)$ for every $x\in[F]$ and $y\in[G]$. Define $z_{j}=\left|f^{-1}\left(j\right)\right|$ and $w_{j}=\left|g^{-1}\left(j\right)\right|$ for $f$ and $g$ and define $z^{\prime}_{j}=\left|(f^{\prime})^{-1}\left(j\right)\right|$ and $w^{\prime}_{j}=\left|(g^{\prime})^{-1}\left(j\right)\right|$ for $f^{\prime}$ and $g^{\prime}$. Then, the following are equivalent:
@@
Proof.  To show that the first item implies the second, let us think of $f^{\prime},g^{\prime}$ as special cases of $(f,g)\in\mathcal{D}$ such that both $f([F])$ and $g([G])$ are contained in $[M^{\prime}]$ (note that such $(f^{\prime},g^{\prime})$ exists in $\mathcal{D}$ since $\mathcal{D}$ is closed under the permutations over the range). In this case, we have $z_{j}=z^{\prime}_{j}$ and $w_{j}=w^{\prime}_{j}$ for all $j\in[M^{\prime}]$, and $z_{i}=w_{j}=0$ for all $j>M^{\prime}$. Thus, it holds that
@@
We next show that the second item implies the first. Fix $(f,g)\in\mathcal{D}$ arbitrarily. Observe then that there exists a permutation $\sigma$ over $[M]$ such that the images of $\sigma f$ and $\sigma g$ are contained in $[M^{\prime}]$, since $\mathrm{dom}(f)$ and $\mathrm{dom}(g)$ have cardinality $F$ and $G$, respectively. Let us define $f^{\prime},g^{\prime}$ by identifying $(f^{\prime},g^{\prime})$ with $(\sigma f,\sigma g)$ as an element in $[M^{\prime}]^{[F]}\times[M^{\prime}]^{[G]}$. Then, it holds that $z^{\prime}_{j}=z_{\sigma^{-1}(j)}$ and $w^{\prime}_{j}=w_{\sigma^{-1}(j)}$ for every $j\in[M^{\prime}]$. By bisymmetry, we have $\Phi(f,g)=\Phi(\sigma f,\sigma g)$, which is equal to $\Phi^{\prime}(f^{\prime},g^{\prime})$ by the definition.
Suppose that the second item holds. Then, $\Phi^{\prime}(f^{\prime},g^{\prime})$ is $\epsilon$-approximated by some polynomial $Q^{\prime}$ of degree at most $d$ over $z^{\prime}_{1},\dots,z^{\prime}_{M^{\prime}},w^{\prime}_{1},\dots,w^{\prime}_{M^{\prime}}$. Since $\Phi^{\prime}$ is a bisymmetric property, the average of $Q^{\prime}$ over all permutations over $[M^{\prime}]$, i.e.,
@@
By combining Sec. missing  3.1 with Sec. missing  3.1, we obtain a slightly general version of Theorem 1.
@@
For $F,G,M\in\mathbb{N}$ such that $M\geq F+G$ and $\mathcal{D}\subseteq[M]^{[F]}\times[M]^{[G]}$, let $\Phi\colon\mathcal{D}\to\{0,1\}$ be a bisymmetric Boolean function over the set of function pairs $(f,g)\in\mathcal{D}$. In addition, for every $M^{\prime}\in\mathbb{N}$ such that $F+G\leq M^{\prime}<M$, let $\Phi^{\prime}\colon\mathcal{D}\cap\left([M^{\prime}]^{[F]}\times[M^{\prime}]^{[G]}\right)\to\{0,1\}$ be the restriction of $\Phi$ to the domain $[M^{\prime}]^{[F]}\times[M^{\prime}]^{[G]}$. Then, the minimum degree of a polynomial that $\epsilon$-approximates $\Phi$ is equal to the minimum degree of a polynomial that $\epsilon$-approximates $\Phi^{\prime}$, where
@@
With the promise that the union of the images of input functions $f$ and $g$, i.e., ${f([F])\cup g([G])}$, is upper-bounded in size by $\mathfrak{U}$ for every $(f,g)\in\mathcal{D}$, Sec. missing  3.1 holds for every $M^{\prime}$ with $\mathfrak{U}\leq M^{\prime}<M$ with a slight modification to proof. Thus, with that promise, Theorem 12 also holds for such $M^{\prime}$. For instance, if $\mathcal{D}$ is the set of the pairs of two-to-one functions, then $\lvert f([F])\cup g([G])\rvert\leq(F+G)/2$, and thus Theorem 12 holds for every $M^{\prime}$ with $(F+G)/2\leq M^{\prime}<M$. Since $\lvert f([F])\cup g([G])\rvert\leq F+G$ always holds, we obtain Theorem 12 for all $\mathcal{D}$.
3.2 $k$-Symmetric Properties
This section states a straightforward generalization of Lemmas 3.1 and 3.1 to the $k$-symmetric properties. We omit their proofs, since their proofs are almost the same as those in the bisymmetric case. Theorem 15 (a slightly general version of Theorem 2) follows immediately from the following lemmas.
@@ 3.2 $k$-Symmetric Properties
Let $k\geq 2$. For $F_{\ell}\in\mathbb{N}\ (\ell\in[k])$, $M\in\mathbb{N}$, let $\Phi:\mathcal{D}\to\{0,1\}$ be a $k$-symmetric partial Boolean function over the set of $k$-tuples of functions $(f_{1},\dots,f_{k})\in\mathcal{D}$, where $\mathcal{D}\subseteq\mathcal{F}_{1}\times\dots\times\mathcal{F}_{k}$ and $\mathcal{F}_{\ell}$ denotes $[M]^{[F_{\ell}]}$ for each $\ell\in[k]$. For each $\ell\in[k]$, $i\in[F_{\ell}]$, and $j\in[M]$, define variable $x^{(\ell)}_{ij}$ such that $x^{(\ell)}_{ij}=1$ if and only if $f_{\ell}\left(i\right)=j$, and define $z^{(\ell)}_{j}$ as the number of preimages of $j\in[M]$ via $f_{\ell}$, i.e., $z^{(\ell)}_{j}=\left|f_{\ell}^{-1}\left(j\right)\right|$. Then, the following are equivalent.
@@
. Let $k\geq 2$. For $F_{\ell}\ (\ell\in[k])$ and $M\in\mathbb{N}$ such that $M\geq F_{1}+\dots+F_{k}$, let $\Phi:\mathcal{D}\to\{0,1\}$ be a $k$-symmetric partial Boolean function over the set of $k$-tuples of functions $(f_{1},\dots,f_{k})\in\mathcal{D}$, where $\mathcal{D}\subseteq\mathcal{F}_{1}\times\dots\times\mathcal{F}_{k}$ and $\mathcal{F}_{\ell}$ denotes $[M]^{[F_{\ell}]}$ for each $\ell\in[k]$. In addition, for every ${M^{\prime}}\in\mathbb{N}$ such that $F_{1}+\dots+F_{k}\leq{M^{\prime}}<M$, let $\Phi^{\prime}:\mathcal{D}\cap\left(\mathcal{F}^{\prime}_{1}\times\dots\times\mathcal{F}^{\prime}_{k}\right)\to\{0,1\}$ be the restriction of $\Phi$ to the domain $[{M^{\prime}}]^{[F_{1}]}\times\dots\times[{M^{\prime}}]^{[F_{k}]}$, i.e., the $k$-symmetric Boolean function over the set of $k$-tuples of functions $(f^{\prime}_{1},\dots,f^{\prime}_{k})\in\mathcal{D}\cap\left(\mathcal{F}^{\prime}_{1}\times\dots\times\mathcal{F}^{\prime}_{k}\right)$, where $\mathcal{F}^{\prime}_{\ell}$ denotes $[{M^{\prime}}]^{[F_{\ell}]}$ for each $\ell\in[k]$, such that $\Phi^{\prime}(f^{\prime}_{1},\dots,f^{\prime}_{k})=\Phi(f_{1},\dots,f_{k})$ for all functions $f^{\prime}_{1},\dots,f^{\prime}_{k},f_{1},\dots,f_{k}$ satisfying $f^{\prime}_{\ell}(x)=f_{\ell}(x)$ for all $x\in[F_{\ell}]$ and all $\ell\in[k]$. Finally, for each $\ell\in[k]$, define $z^{(\ell)}_{j}=\left|f_{\ell}^{-1}\left(j\right)\right|$ for each $j\in[M]$ and ${z^{\prime}}^{(\ell)}_{j}=\left|(f^{\prime}_{\ell})^{-1}\left(j\right)\right|$ for each $j\in[{M^{\prime}}]$. Then, the following are equivalent:
@@
Define $\Phi$ and $\Phi^{\prime}$ as in Sec. missing  3.2, where
@@
Then, the minimum degree of a polynomial that $\epsilon$-approximates $\Phi$ is equal to the minimum degree of a polynomial that $\epsilon$-approximates $\Phi^{\prime}$.

### v2

@@ 2. $i$
If $h$ is one-to-one, then $h|_{S}$ and $h|_{T}$ obviously have no collision for every $S$ and $T$, where $h|_{S}$ and $h|_{T}$ are the restriction of $h$ to the the subsets $S$ and $T$, respectively, of the domain. Thus, $P$ is close to 0. If $h$ is two-to-one, $h|_{S}$ is injective with high probability and, in this case, $T$ intersects the preimage of $h|_{S}(S)$ with high probability. This follows from the birthday paradox, since $|S|=F<<\sqrt{M}=\sqrt{FG}$ and $|T|=G>>\sqrt{M}$. Therefore, $P$ is far from 0. By scaling and shifting $P$, we have a polynomial that $O(1)$-approximates the collision property.
As an application of Theorem 4, we obtain Theorem 5 by generalizing the reduction from the claw problem for two functions with the same domain size [7] to the reduction from the problem for two functions with the different domain size: $f\colon[F]\to[F+G]$ and $g\colon[G]\to[F+G]$ with $F\leq G\leq F^{2}$ (the case of $G>F^{2}$ is already covered by Theorem 4). More precisely, we reduce the composition of a certain problem (called $\mathsf{pSearch}$ [11]) and the claw problem with the function range stated in Theorem 4 to the claw problem in question, and then use the composition theorem of the general adversary bound [19]. That is, for every $k,F,G\in\mathbb{N}$ with $k|F$ and $G|(F/k)$, we show the composition $\mathsf{claw}_{(F,G)\to M+2}$ is reducible from the problem $\mathsf{claw}_{(k,kG/F)\to M}\circ\mathsf{pSearch}_{F/k\to M}$, where $\mathsf{claw}_{(a,b)\to c}$ denotes the decision problem of detecting the existence of a claw for given functions $(\phi,\psi)\in[c]^{[a]}\times[c]^{[b]}$. Since $\mathsf{pSearch}_{N\to M}$ is an variant of the unstructured search over $N$ elements, the quantum query complexity of $\mathsf{pSearch}_{F/k\to M}$ is $\Omega(\sqrt{F/k})$ [11]. By setting $M=k+kG/F$, Theorem 4 provides a lower bound on the quantum query complexity of $\mathsf{claw}_{(k,kG/F)\to M}$ as
@@
By applying the composition theorem of the general adversary bound, we have
@@ 1.5 Organization
Sec.   2 provides notations and necessary notions. Sec.   3 proves Theorem 1 and Theorem 2. Sec.   4 proves Theorem 3 and Theorem 4. Sec.   5 proves Theorem 5.
@@ 2.1 Notations
For any ordered sets $\mathcal{D}$ and $\mathcal{R}$, we identify a function $f\colon\mathcal{D}\to\mathcal{R}$ as the sequence of $(f(d_{1}),\dots,f(d_{|\mathcal{D}|}))$, where $d_{k}$ is the $k$th element in $\mathcal{D}$ for each $k\in[|\mathcal{D}|]$. We thus represent by $\mathcal{R}^{\mathcal{D}}$ the set of functions from $\mathcal{D}$ to $\mathcal{R}$. We may abbreviate $\mathcal{R}^{[D]}$ as $\mathcal{R}^{D}$, when $\mathcal{D}=[D]$. For instance, $M]^{[F]}$ (or $M]^{F}$) for $F,M\in\mathbb{N}$ denotes the set of functions from $\left[F\right]$ to $\left[M\right]$.
@@ 2.2 Polynomial Method for Quantum Lower Bound
The polynomial method was introduced by Beals et al. [9] to obtain lower bounds on the number of quantum queries required to compute a Boolean function on an input bit-string given as an oracle. It is not difficult to generalize the method to the case where the input is a string of alphabets [1]. In the following, we consider the case where the input consists of two alphabet strings corresponding to two functions. However, it is not difficult to see similar facts hold for three or more functions.
The following lemma says that, after $q$ quantum queries, any quantum algorithm is in a state where the amplitude of each computational basis state is represented by a polynomial of degree at most $q$ in $x_{ij}$ and $y_{kj}$.
@@
Let $F,G,M$ be in $\mathbb{N}$. For given oracles $O_{f}$ and $O_{g}$ for functions $f\in[M]^{[F]}$ and $g\in[M]^{[G]}$, assume, without loss of generality, that any quantum algorithm with $q$ queries applies
@@
where ++ is the addition modulo $M$, or alternatively,
@@
where $\oplus$ is bitwise XOR in binary expression. At any step, the state of the algorithm is of the form
@@
Let $F,G,M$ be in $\mathbb{N}$, and let $\Phi$ be a partial Boolean function over the set of pairs of functions in $\mathcal{D}\subseteq[M]^{[F]}\times[M]^{[G]}$. For given oracles $O_{f}$ and $O_{g}$ for functions $(f,g)\in\mathcal{D}$, suppose that there exists a quantum algorithm that computes $\Phi(f,g)$ with error $\epsilon$ using $q$ queries. Then, there is a polynomial $P$ of degree at most $2q$ in variables $x_{11},\dots,x_{FM},y_{11},\dots,y_{GM}$ such that $P$ $\epsilon$-approximates $\Phi$, where $x_{ij}$ and $y_{kj}$ are the variables defined for $f$ and $g$, respectively, in Sec.   2.2.
@@
For a partial Boolean function $\Phi$ and $0<\epsilon<\delta<1/2$, it holds that
@@ 2.3 Multisymmetric Properties
The definition of bisymmetric properties can be generalized to the three or more function case in a straightforward way. Let $\Phi\colon\mathcal{D}\to\{0,1\}$ be a partial Boolean function over the set of $k$-tuples of functions in $\mathcal{D}\subseteq\mathcal{F}_{1}\times\cdots\times\mathcal{F}_{k}$, where $\mathcal{F}_{\ell}\ (\ell\in[k])$ denotes $M]^{[F_{\ell}]}$ for $F_{\ell},M\in\mathbb{Z}_{+}$. We say that $\Phi$ is $k$-symmetric, if $\Phi(f_{1},\dots,f_{k})$ is invariant and $\mathcal{D}$ is closed under the permutations over $\mathrm{dom}(f_{\ell})=[F_{\ell}]$ for each $\ell\in[k]$, respectively, and the permutation over the common range $[M]$ of $f_{\ell}$’s, i.e., $\left(\sigma f_{1}\pi_{F_{1}},\dots,\sigma f_{k}\pi_{F_{k}}\right)\in\mathcal{D}$ if $(f_{1},\dots,f_{k})\in\mathcal{D}$, and
@@
for every permutation $\pi_{F_{\ell}}$ over $F_{\ell}]$ for $\ell\in[k]$ and every permutation $\sigma$ over $[M]$, respectively.
@@ 3.1 Bisymmetric Properties
We first show that there are two polynomials in different sets of variables of the same degree such that both polynomials approximate $\Phi$.
@@
For $F,G,M\in\mathbb{N}$ and $\mathcal{D}\subseteq[M]^{[F]}\times[M]^{[G]}$, let $\Phi:\mathcal{D}\to\{0,1\}$ be a bisymmetric partial Boolean function over the set of functions $(f,g)\in\mathcal{D}$. Define $FM$ variables $x_{ij}$ and $GM$ variables $y_{kj}$ for functions $f$ and $g$ as in Sec.   2.2. In addition, define $z_{j}$ as the number of preimages of $j$ via $f$, i.e., $z_{j}=\left|f^{-1}\left(j\right)\right|$, and define $w_{j}$ as the number of preimages of $j$ via $g$, i.e., $w_{j}=\left|g^{-1}\left(j\right)\right|$. Then, the following are equivalent.
@@
Next, we show that the degree of a polynomial $Q$ shown in Sec.   3.1 does not change, even if we restrict the common range of input functions of $\Phi$ to a small subset.
@@
For $F,G,M\in\mathbb{N}$ such that $M\geq F+G$ and $\mathcal{D}\subseteq[M]^{[F]}\times[M]^{[G]}$, let $\Phi\colon\mathcal{D}\to\{0,1\}$ be a bisymmetric partial Boolean function over the set of function pairs $(f,g)\in\mathcal{D}$. In addition, for every $M^{\prime}\in\mathbb{N}$ such that $F+G\leq M^{\prime}<M$, let $\Phi^{\prime}\colon\mathcal{D}\cap\left([M^{\prime}]^{[F]}\times[M^{\prime}]^{[G]}\right)\to\{0,1\}$ be the restriction of $\Phi$ to the domain $M^{\prime}]^{[F]}\times[M^{\prime}]^{[G]}$, i.e., the bisymmetric Boolean function over the set of function pairs $(f^{\prime},g^{\prime})\in\mathcal{D}\cap\left([M^{\prime}]^{[F]}\times[M^{\prime}]^{[G]}\right)$ such that $\Phi^{\prime}(f^{\prime},g^{\prime})=\Phi(f,g)$ for all functions $f^{\prime},g^{\prime},f,g$ such that $f^{\prime}(x)=f(x)$ and $g^{\prime}(y)=g(y)$ for every $x\in[F]$ and $y\in[G]$. Define $z_{j}=\left|f^{-1}\left(j\right)\right|$ and $w_{j}=\left|g^{-1}\left(j\right)\right|$ for $f$ and $g$ and define $z^{\prime}_{j}=\left|(f^{\prime})^{-1}\left(j\right)\right|$ and $w^{\prime}_{j}=\left|(g^{\prime})^{-1}\left(j\right)\right|$ for $f^{\prime}$ and $g^{\prime}$. Then, the following are equivalent:
@@
Proof. To show that the first item implies the second, let us think of $f^{\prime},g^{\prime}$ as special cases of $(f,g)\in\mathcal{D}$ such that both $f([F])$ and $g([G])$ are contained in $M^{\prime}]$ (note that such $(f^{\prime},g^{\prime})$ exists in $\mathcal{D}$ since $\mathcal{D}$ is closed under the permutations over the range). In this case, we have $z_{j}=z^{\prime}_{j}$ and $w_{j}=w^{\prime}_{j}$ for all $j\in[M^{\prime}]$, and $z_{i}=w_{j}=0$ for all $j>M^{\prime}$. Thus, it holds that
@@
We next show that the second item implies the first. Fix $(f,g)\in\mathcal{D}$ arbitrarily. Observe then that there exists a permutation $\sigma$ over $[M]$ such that the images of $\sigma f$ and $\sigma g$ are contained in $M^{\prime}]$, since $\mathrm{dom}(f)$ and $\mathrm{dom}(g)$ have cardinality $F$ and $G$, respectively. Let us define $f^{\prime},g^{\prime}$ by identifying $(f^{\prime},g^{\prime})$ with $(\sigma f,\sigma g)$ as an element in $M^{\prime}]^{[F]}\times[M^{\prime}]^{[G]}$. Then, it holds that $z^{\prime}_{j}=z_{\sigma^{-1}(j)}$ and $w^{\prime}_{j}=w_{\sigma^{-1}(j)}$ for every $j\in[M^{\prime}]$. By bisymmetry, we have $\Phi(f,g)=\Phi(\sigma f,\sigma g)$, which is equal to $\Phi^{\prime}(f^{\prime},g^{\prime})$ by the definition.
Suppose that the second item holds. Then, $\Phi^{\prime}(f^{\prime},g^{\prime})$ is $\epsilon$-approximated by some polynomial $Q^{\prime}$ of degree at most $d$ over $z^{\prime}_{1},\dots,z^{\prime}_{M^{\prime}},w^{\prime}_{1},\dots,w^{\prime}_{M^{\prime}}$. Since $\Phi^{\prime}$ is a bisymmetric property, the average of $Q^{\prime}$ over all permutations over $M^{\prime}]$, i.e.,
@@
By combining Sec.   3.1 with Sec.   3.1, we obtain a slightly general version of Theorem 1.
@@
For $F,G,M\in\mathbb{N}$ such that $M\geq F+G$ and $\mathcal{D}\subseteq[M]^{[F]}\times[M]^{[G]}$, let $\Phi\colon\mathcal{D}\to\{0,1\}$ be a bisymmetric Boolean function over the set of function pairs $(f,g)\in\mathcal{D}$. In addition, for every $M^{\prime}\in\mathbb{N}$ such that $F+G\leq M^{\prime}<M$, let $\Phi^{\prime}\colon\mathcal{D}\cap\left([M^{\prime}]^{[F]}\times[M^{\prime}]^{[G]}\right)\to\{0,1\}$ be the restriction of $\Phi$ to the domain $M^{\prime}]^{[F]}\times[M^{\prime}]^{[G]}$. Then, the minimum degree of a polynomial that $\epsilon$-approximates $\Phi$ is equal to the minimum degree of a polynomial that $\epsilon$-approximates $\Phi^{\prime}$, where
@@
With the promise that the union of the images of input functions $f$ and $g$, i.e., $f([F])\cup g([G])$}, is upper-bounded in size by $\mathfrak{U}$ for every $(f,g)\in\mathcal{D}$, Sec.   3.1 holds for every $M^{\prime}$ with $\mathfrak{U}\leq M^{\prime}<M$ with a slight modification to the proof. Thus, with that promise, Theorem 12 also holds for such $M^{\prime}$. For instance, if $\mathcal{D}$ is the set of pairs of two-to-one functions, then $\lvert f([F])\cup g([G])\rvert\leq(F+G)/2$, and thus Theorem 12 holds for every $M^{\prime}$ with $(F+G)/2\leq M^{\prime}<M$. Since $\lvert f([F])\cup g([G])\rvert\leq F+G$ always holds, we obtain Theorem 12 for all $\mathcal{D}$.
[…]
This section states a straightforward generalization of Lemmas 3.1 and 3.1 to the $k$-symmetric properties. We omit their proofs, since they are almost the same as those in the bisymmetric case. Theorem 15 (a slightly general version of Theorem 2) follows immediately from the following lemmas.
@@
Let $k\geq 2$. For $F_{\ell}\in\mathbb{N}\ (\ell\in[k])$, $M\in\mathbb{N}$, let $\Phi:\mathcal{D}\to\{0,1\}$ be a $k$-symmetric partial Boolean function over the set of $k$-tuples of functions $(f_{1},\dots,f_{k})\in\mathcal{D}$, where $\mathcal{D}\subseteq\mathcal{F}_{1}\times\dots\times\mathcal{F}_{k}$ and $\mathcal{F}_{\ell}$ denotes $M]^{[F_{\ell}]}$ for each $\ell\in[k]$. For each $\ell\in[k]$, $i\in[F_{\ell}]$, and $j\in[M]$, define variable $x^{(\ell)}_{ij}$ such that $x^{(\ell)}_{ij}=1$ if and only if $f_{\ell}\left(i\right)=j$, and define $z^{(\ell)}_{j}$ as the number of preimages of $j\in[M]$ via $f_{\ell}$, i.e., $z^{(\ell)}_{j}=\left|f_{\ell}^{-1}\left(j\right)\right|$. Then, the following are equivalent.
@@
. Let $k\geq 2$. For $F_{\ell}\ (\ell\in[k])$ and $M\in\mathbb{N}$ such that $M\geq F_{1}+\dots+F_{k}$, let $\Phi:\mathcal{D}\to\{0,1\}$ be a $k$-symmetric partial Boolean function over the set of $k$-tuples of functions $(f_{1},\dots,f_{k})\in\mathcal{D}$, where $\mathcal{D}\subseteq\mathcal{F}_{1}\times\dots\times\mathcal{F}_{k}$ and $\mathcal{F}_{\ell}$ denotes $M]^{[F_{\ell}]}$ for each $\ell\in[k]$. In addition, for every $M^{\prime}$$\in\mathbb{N}$ such that $F_{1}+\dots+F_{k}\leq{M^{\prime}}<M$, let $\Phi^{\prime}:\mathcal{D}\cap\left(\mathcal{F}^{\prime}_{1}\times\dots\times\mathcal{F}^{\prime}_{k}\right)\to\{0,1\}$ be the restriction of $\Phi$ to the domain ${M^{\prime}}]^{[F_{1}]}\times\dots\times[{M^{\prime}}]^{[F_{k}]}$, i.e., the $k$-symmetric Boolean function over the set of $k$-tuples of functions $(f^{\prime}_{1},\dots,f^{\prime}_{k})\in\mathcal{D}\cap\left(\mathcal{F}^{\prime}_{1}\times\dots\times\mathcal{F}^{\prime}_{k}\right)$, where $\mathcal{F}^{\prime}_{\ell}$ denotes ${M^{\prime}}]^{[F_{\ell}]}$ for each $\ell\in[k]$, such that $\Phi^{\prime}(f^{\prime}_{1},\dots,f^{\prime}_{k})=\Phi(f_{1},\dots,f_{k})$ for all functions $f^{\prime}_{1},\dots,f^{\prime}_{k},f_{1},\dots,f_{k}$ satisfying $f^{\prime}_{\ell}(x)=f_{\ell}(x)$ for all $x\in[F_{\ell}]$ and all $\ell\in[k]$. Finally, for each $\ell\in[k]$, define $z^{(\ell)}_{j}=\left|f_{\ell}^{-1}\left(j\right)\right|$ for each $j\in[M]$ and ${z^{\prime}}^{(\ell)}_{j}=\left|(f^{\prime}_{\ell})^{-1}\left(j\right)\right|$ for each $j\in[{M^{\prime}}]$. Then, the following are equivalent:
@@
Define $\Phi$ and $\Phi^{\prime}$ as in Sec.   3.2, where
@@
Then, the minimum degree of a polynomial that $\epsilon$-approximates $\Phi$ equals the minimum degree of a polynomial that $\epsilon$-approximates $\Phi^{\prime}$.